A car accelerates uniformly from 5m/s to 13 m/s in 4 seconds. What is the acceleration of the car in m/s-2
1
Expert's answer
2020-12-04T09:35:50-0500

We can find the acceleration of the car from the kinematic equation:


"v=v_0+at,""a=\dfrac{v-v_0}{t}=\dfrac{13\ \dfrac{m}{s}-5\ \dfrac{m}{s}}{4\ s}=2\ \dfrac{m}{s^2}."

Answer:

"a=2\ \dfrac{m}{s^2}."